# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v4

@@ 4.3 Antifragility and the antifragile terms in RL
(a) $H(t)$ for the first difference $h_{i}(t)$.
(b) $\Delta H(t)$ for the second difference $\Delta h_{i}(t)$.
Figure 4: Illustration of the term $H(t)$ and its effect on the MFD
The second derivative $\Delta H(t)$ is shown in Fig. 5. The y-axis represents how fast the traffic state changes, and the faster it reaches $n_{i,\rm crit}$, the greater the penalty will be. This observation is consistent with the redundant overcompensation and time-scale separation principles formalized in Taleb and Douady (2013) and practically applied in Axenie and Saveriano (2023). On the contrary, a reward will be given if $n$ decelerates when approaching $n_{i,\rm crit}$. Likewise, $\Delta H(t)$ is also dependent on the normalization factor $\omega_{\Delta h}$. With $H(t)$ and $\Delta H(t)$, the agent learns to be conservative when regulating the perimeter control variables to reach $n_{i,\rm crit}$.
Figure 5: Illustration of the term $\Delta H(t)$
@@ 5 Experiment application
    Baseline RL perimeter control algorithm proposed in Zhou and Gayah (2021).
It is worth mentioning that this work employs MPC-MHE differently in comparison to Sirmatel and Geroliminis (2020), where sensor measurement noise is assumed and addressed by MHE through state estimation. In contrast, we presume the presence of disruption but perfect sensor measurements as the real states. As a result, since there is no need to estimate the system states, MPC-MHE does not provide additional benefits over MPC alone in the context of demand disruptions. However, for supply disruptions, we introduce a disruption magnitude reduction coefficient as an additional state variable, and MPC-MHE can be applied to estimate this coefficient and thereby the shift of the model.
Since transportation systems inherently manage the imbalance between demand and supply, real-world disruptions can also be broadly categorized as either a demand or a supply disruption. Therefore, the performance of the studied perimeter control algorithms in this paper is tested under demand or supply disruptions, as illustrated in Fig. 6. A demand disruption can be easily understood as a surging traffic demand due to social events or similar occurrences. Population growth and urbanization can also be considered a long-term and gradually incremental demand disruption. On the other side, a supply disruption can represent a link-level or network-level capacity drop due to adverse weather or major accidents, with network-level capacity drop being able to be reflected by the decrease of the MFD profile, as in Lu et al. (2024). Increasing traffic demand often leads to more frequent and severe traffic accidents (Dickerson et al., 2000). Other than the deterministically linearly increasing magnitude of disruptions, as represented by the solid linear red line, we also tested the algorithms of incremental magnitudes with uncertainties taken into account, as illustrated by the zigzag dotted red line, which will be explained in detail in the following section.
Figure 6: Illustration of simulation scenarios and their evaluation
@@ 5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 2(b) shows. These MFDs are largely the same as in Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, the lack of differentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by merely $3\%$, so that the whole MFD can be both continuous and differentiable. Other important indicators, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
@@
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 7(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.2,0.4,0.1,0.3]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,N}(t)=C_{ij,N}\mathcal{N}_{ij}(\mu_{ij,N},\sigma_{ij,N})$ with total number of vehicles following Gaussian as $C_{ij,N}=[3000,10000,2000,7000]$ and the mean and variance of the distribution being $\mu_{ij,N}=[1800,1800,1800,1800]$ and $\sigma_{ij,N}=[1200,1500,900,1200]$ in seconds. Even though real-world disruptions can happen to both the outer and the inner regions, exerting a negative influence on the overall performance, we focus on the critical scenarios where the potential of perimeter control can be better reflected. As the inner region has a significantly smaller MFD profile compared to the outer region, hence, a surging disruptive traffic demand is assumed to be generated within the inner region. Therefore, when testing the performance of different algorithms under a growing demand disruption, $q_{22,N}$ is increased to simulate traffic from within the city center, reaching a total disruption demand of $12000$ vehicles at episode $75$, accounting for approximately one-third of the normal demand.
(a) Base demand profile and demand disruption
(b) MFD and supply disruption for the inner region
Figure 7: Demand profile with or without surging demand
@@
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. In this work, we adopt a proportional decrease in both jam density and maximal completion of the MFD, as shown in Fig. 7(b), in resemblance to Kim and Yeo (2017); Lu et al. (2024); Sun et al. (2024), representing lane closure events due to road construction works, traffic accidents, and natural disasters. Similar to testing the critical demand disruption on $q_{22,N}$, the critical scenario for supply disruption is the inner MFD being compromised while the outer MFD remains intact. Note that supply disruptions can be modeled in different ways for other events. For instance, considering the performance loss due to vehicle infrastructure interaction, Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is utilized, and a smaller $\lambda$ indicates the infrastructure is used with higher efficiency.
@@
The total simulation duration $T$ is 3 hours, with each time step $\Delta t$ taking $180$ seconds, and the third and last hour has little demand as it mainly serves as the unloading process. Another constraint is the lower and upper bounds for the perimeter control variable $u_{ij}(t)\in[0.1,0.9]$ to represent the real-world constraints from traffic signals (Geroliminis et al., 2013). The initial vehicle accumulation is set to be $n_{ij}(0)=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation. Each scenario is run $25$ times since randomness is inherent in RL, and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 6. A list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The same multiplier list is employed for all scenarios that are subject to disruption magnitude uncertainties. Also, it is shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through reshuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers but with different sequences.
@@
The most important hyperparameters for both the baseline and the proposed RL algorithms should be the same and are summarized in Table 2. Note that the minimal learning rates and noise scale are not set to be a rather small value as common RL algorithms do, we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions, which can be demonstrated by the results for both superior performance and antifragility in Section 6. The coefficients $\xi_{1}$ and $\xi_{2}$ for the damping term $r_{dam}(t)$ in Eq. 10 to counter the potential oscillations actions are set to be $1$ and $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red}(t)$ is set to be $0.1$ and $0.2$. The number of simulations generated under the Ape-X architecture is 32 per training episode. Both the prediction horizon and control horizon of MPC and MPC-MHE are set to 10.
Table 2: List of hyperparameters
@@ 5.2 Performance evaluation
The reward in the RL algorithm is defined based on trip completion with additional damping and redundancy terms. However, two scenarios with the same trip completion at the end of the simulation may exhibit distinct Total Time Spent (TTS), and the one with the lower value should be regarded as having demonstrated a superior overall performance. In addition, to better showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 16, the main performance indicator being evaluated in this work is the TTS, which is computed by adding up the number of vehicles within the network at each second of the simulation.
Since urban road networks are always subject to capacity constraints, and as Sun et al. (2024) has recently proven the fragile nature of urban road networks, fully antifragile traffic control strategies may be impossible to achieve and will inevitably fall into being fragile when the magnitude of disruptions is large enough. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is less fragile than the state-of-the-art baseline algorithms, and we normalize all the other perimeter control algorithms over the RL baseline method to study the relative antifragility. To quantify antifragility of different algorithms, we calculate the distribution skewness based on the samples from the last 25 incremental episodes, with $\mu$ and $\sigma$ denoting the mean and the standard deviation:
@@
A skewness computed to be 00 means that the system itself or the applied algorithm makes the system neither fragile nor antifragile. On the other side, negative skewness indicates the distribution has a longer or fatter right tail and thus a higher degree of concavity in the performance function, which showcases antifragility, whereas a positive skewness signifies fragility. For the detailed methodology regarding skewness and antifragility, interested readers may refer to Sun et al. (2024). Therefore, other than demonstrating a superior performance regarding a lower TTS, the proposed method should also showcase a lower skewness than the benchmarks.
@@ 6.1 Performance under full observability
Under idealized full observability, all algorithms are given perfect information on $[n_{ij}(t),q_{ij}(t)]$. As discussed in Section 5, MPC-MHE is only applied under supply disruption, where a linearly increasing reduction factor $r$ is assumed and can be estimated through MHE.
@@ 6.1.1 Demand disruption
Demand disruption can be understood as a surging traffic demand within a short period, as illustrated in Fig. 7(a) on top of $q_{22,N}$ as a critical scenario. With Fig. 8(a) we present the performance curves of different studied algorithms, i.e., MPC as in Geroliminis et al. (2013) in orange, baseline RL in Zhou and Gayah (2021) in gray, and the proposed antifragile RL algorithm in blue under the cordon-shaped perimeter control environment. The first 50 episodes represent the training process for the RL-based algorithms under a static demand profile, which follows a normal distribution. The TTS curves first drop down and reach a comparable performance as MPC, corresponding to the conclusion in Zhou and Gayah (2021). The 50 training episodes are followed by another 25 episodes with linearly increasing demand disruption, with no stochasticity assumed on top of the linearly growing magnitude. The studied algorithms showcase distinct capabilities of learning and adapting to such disruptions, and the performance variance from the 25 iterations of the simulation. It is obvious that the proposed antifragile RL algorithm exhibits both superior performance and reduced variance, as indicated by the significantly narrower shaded area of its curve. Furthermore, the performance curve of the proposed method also seems less convex than the other algorithms, demonstrating its relative antifragility. Note that due to the inherent variability of RL-based algorithms as well as the high sensitivity of the distribution skewness, which will be computed below, on such variability, we apply a sliding window of 5 episodes to smooth the performance curve.
@@
Such superior performance is further quantitatively illustrated with a polar plot in Fig. 8(b), which shows the performance difference in TTS reduction normalized over the baseline RL algorithm. Only the episodes $50-75$ under incremental demand are shown. Instead of variance, the shaded blue area highlights the performance gain of our proposed method compared to the baseline RL algorithm, showing a steady increase in TTS reduction and being largely concave. The average performance gain is $9.3\%$reaching $21.0\%$ after 25 episodes under incremental demand disruption. On the other hand, although MPC performs relatively better when the magnitude of disruption is low, its performance quickly draws near to the baseline RL under high demand disruptions, achieving only a $4.8\%$ improvement by the end of the simulation.
@@
To quantify the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. 8(c), meaning the value of the skewness curve at the final episode reflects the antifragility characteristics across all $25$ episodes under incremental disruptions. Since the calculation of skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase, i.e., episode $50-55$. Initially, the skewness for all three methods starts at a low value. However, the skewness of the baseline RL in gray rises rapidly, whereas the skewness of the MPC algorithm progresses even faster to $0.86$, despite the good performance when the magnitude of disruption is low, as depicted in Fig. 8(c). The skewness of the antifragile RL-based algorithm in blue, on the contrary, increases more slowly, and lies largely below the baseline curve, especially when the magnitude of disruption is high, ultimately achieving the lowest final skewness at $0.46$.
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 8: Performance curves under incremental demand disruptions
(a) Performance curves in TTS
(b) Normalized over baseline RL
(c) Skewness up to the $n$-th episode
Figure 9: Performance curves under incremental demand disruptions with uncertainty
@@
Fig. 9 shows the performance when stochasticity is introduced to the magnitude of disruptions, following a disruption shuffling scheme discussed in Section 5.1 to ensure each episode between $50-75$ goes through the same total magnitude of disruptions. The results display a consistent trend with the non-stochastic scenario, with the performance curves being less smooth, as in Fig. 9(b), suggesting that the presence of uncertainty adds complexity to the training process. The performance of both RL-based methods deteriorates slightly, but still, the performance of the proposed antifragile RL algorithm outperforms the baseline RL by an average of $8.0\%$ and achieves $18.6\%$ performance gain at the end of the simulation under incremental demand disruptions, as well as a gain of $12.9\%$ compared to MPC. The fact that the TTS difference curve is approximately convex for MPC and concave for the antifragile RL algorithm can be quantified in Fig. 9(c). While the final distribution skewness of the baseline and the antifragile RL algorithms is $0.70$ and $0.56$, respectively, the skewness of MPC records a value of $0.93$, indicating its high fragility under incremental demand disruptions. An interesting phenomenon worth noticing is the spike of the skewness curve of the proposed algorithm around episode $60-65$, which can be observed in not only Fig. 8(c) and 9(c), but also in the results from supply disruption in the following discussion. The observed effect is likely due to the interplay between multiple RL hyperparameters, including the memory buffer, step size, and learning rate. After episode $50$, the replay buffer contains a mixture of training data under both disrupted and undisrupted scenarios, resulting in learning from a mixed dataset and causing a moderate deviation from optimal performance.

### v5

@@ 4.3 Antifragility and the antifragile terms in RL
(a) $H_{k}$ for the first difference $h_{i,k}$.
(b) The impact on the MFD with $n_{i,k}\geq n_{i,k-1}$.
Figure 4: Illustration of the term $H_{k}$ and its effect on the MFD.
The second derivative $\Delta H_{k}$ is shown in Fig. 5. The y-axis represents how fast the traffic state changes, and the faster it reaches $n_{i,\rm crit}$, the greater the penalty will be. This observation is consistent with the redundant overcompensation and time-scale separation principles formalized in Taleb and Douady (2013) and practically applied in Axenie and Saveriano (2023). On the contrary, a reward will be given if $n$ decelerates when approaching $n_{i,\rm crit}$. Likewise, $\Delta H_{k}$ is also dependent on the normalization factor $\omega_{\Delta h}$. With $H_{k}$ and $\Delta H_{k}$, the agent learns to be conservative when regulating the perimeter control variables to reach $n_{i,\rm crit}$.
Figure 5: Illustration of the term $\Delta H_{k}$.
@@ 5 Experiment application
    Baseline RL perimeter control algorithm proposed in Zhou and Gayah (2021) with the original codebase [^1] https://github.com/DongqinZhou/C-D-RL.
It is worth mentioning that this work employs MPC-MHE differently in comparison to Sirmatel and Geroliminis (2020), where sensor measurement noise is assumed and addressed by MHE through state estimation. In contrast, we presume the presence of disruption but perfect sensor measurements as the real states. As a result, since there is no need to estimate the system states, MPC-MHE does not provide additional benefits over MPC alone in the context of demand disruptions. For supply disruptions, however, we introduce a disruption magnitude reduction coefficient as an additional state variable, and MPC-MHE can help estimate the coefficient and thereby the model shift, supposing such variation is well-defined and can be modeled.
Since transportation systems inherently manage the imbalance between demand and supply, real-world disruptions can also be broadly categorized as either a demand or supply disruption. A demand disruption can be easily understood as a surging traffic demand due to social events or other occurrences. Population growth and urbanization can also be considered a long-term and gradually incremental demand disruption. On the other hand, a supply disruption can represent a link-level or network-level capacity drop due to adverse weather or major accidents, with the network capacity drop being reflected by the decrease of the MFD profile, as in Lu et al. (2024). In this paper, the performance of the studied perimeter control algorithms is also tested under demand or supply disruptions, as illustrated in Fig. 6. Increasing traffic demand often leads to more frequent and severe traffic accidents (Dickerson et al., 2000). In addition to the deterministically linearly increasing magnitude of disruptions, as represented by the solid linear red line, we also tested the algorithms with uncertainties taken into account, as illustrated by the zigzag dotted red line, which will be explained in detail in the following section.
Figure 6: Illustration of simulation scenarios and their evaluation.
@@ 5.1 Simulation environment parametrization
We simulate a cordon-shaped urban network with inner and outer regions represented by different MFDs, as Fig. 2(b) shows. These MFDs follow Zhou and Gayah (2021) and were originally approximated through the Yokohama loop detector dataset (Geroliminis and Daganzo, 2008). However, since the MFDs in Zhou and Gayah (2021) are formulated as piecewise functions and are, although continuous, not differentiable, the lack of differentiability within MFDs can cause fluctuations when computing the first and second differences and harm the efficacy of the redundancy term. Therefore, a minor modification has been made to slightly increase the gridlock accumulation by merely $3\%$, so that the whole MFD can be both continuous and differentiable. Other essential parameters, e.g., critical accumulation and maximal trip completion, remain the same as in Zhou and Gayah (2021).
@@
Traffic demand under no disruption is approximated based on Geroliminis and Daganzo (2008). As the real-world peak hour traffic demand profile has more resemblance to a Gaussian distribution instead of being a simple trapezoidal shape (Mazloumi et al., 2010), the base demand $q_{ij}(t)=[q_{11}(t),q_{12}(t),q_{21}(t),q_{22}(t)]$ is illustrated as the blue curves in Fig. 7(a). It consists of two components, which are a constant value $q_{ij,c}=[q_{11,c},q_{12,c},q_{21,c},q_{22,c}]=[0.2,0.4,0.1,0.3]$ in $\mathrm{veh/s}$ and a Gaussian term $q_{ij,G}(t)=C_{ij}\mathcal{N}_{ij}(\mu_{ij},\sigma_{ij})$ with total number of vehicles following Gaussian as $C_{ij}=[3000,10000,2000,7000]$ and the mean and variance of the distribution being $\mu_{ij}=[1800,1800,1800,1800]$ and $\sigma_{ij}=[1200,1500,900,1200]$ in seconds. Although real-world disruptions can occur in both the outer and inner regions, exerting a negative impact on overall performance, we focus on the critical scenarios where the potential of perimeter control can be better reflected. Since the inner region has a significantly smaller MFD profile compared to its outer counterpart, a surging disruptive traffic demand is presumed to be generated in the inner region. Numerically, $q_{22,G}$ is increased to simulate traffic from within the city center, reaching a total disruption demand of $12000$ vehicles at episode $75$, accounting for approximately one-third of the normal demand.
(a) Base demand profile and demand disruption.
(b) MFD and supply disruption for the inner region.
Figure 7: Demand profile with or without surging demand.
@@
Supply disruptions can manifest in various forms, such as adverse weather conditions, traffic accidents, and road maintenance. However, limited research has been dedicated to understanding the exact correlation between such events and their impact on the MFD. In this work, we adopt a proportional decrease in both maximal accumulation and completion of the MFD, as shown in Fig. 7(b), in resemblance to Lu et al. (2024), representing lane closure events due to road construction works, traffic accidents, and natural disasters. Similar to the critical demand disruption on $q_{22,G}$, critical supply disruption requires the inner MFD being compromised while the outer MFD remains intact. Note that supply disruptions can be modeled in different ways for other events. For instance, considering the performance loss due to vehicle infrastructure interaction, Ambühl et al. (2020) proposed the concept of infrastructure potential $\lambda$ to represent how efficiently the network infrastructure is utilized, with a smaller $\lambda$ indicating the infrastructure is operated with higher efficiency.
@@
The total simulation duration $T$ is 3 hours, and each time step $\Delta t$ takes $180$ seconds. The third and last hour has little demand and mainly serves as the unloading process. Another constraint is the lower and upper bounds for the perimeter control variable $u_{ij,k}\in[0.1,0.9]$ as explained in Eq. 3c. The initial vehicle accumulation is set to be $n_{ij,0}=[600,1300,300,2400]$ so that the accumulation remains approximately at an equilibrium at the beginning of the simulation. Each scenario is run $25$ times since randomness is inherent in RL and its performance may vary across simulation iterations. Each iteration lasts for $75$ episodes, with the first $50$ under no additional disruption so that the RL agent can be properly trained under the base demand profile with predefined training noise. Incremental magnitude of disruptions is introduced in the subsequent $25$ episodes, with the value deliberately set to be the same as the number of simulation iterations. Under disruption uncertainties, as illustrated by the zigzag lines for the simulation scenarios in Fig. 6, a list of multipliers $\epsilon=[\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{24},\epsilon_{25}]$ following a normal distribution $\mathcal{N}\sim(1,0.15)$ is randomly generated to introduce uncertainties by computing the dot product with the list of disruption magnitudes. The same multiplier list is applied for all scenarios that are subject to disruption magnitude uncertainties. Also, it is shuffled by 1 for each simulation iteration. For example, the list would be $\epsilon=[\epsilon_{2},\epsilon_{3},\ldots,\epsilon_{25},\epsilon_{1}]$ for the second iteration. Through shuffling, the disruption magnitudes from different episodes will experience exactly the same list of uncertainty multipliers, but with different sequences.
@@
The most important hyperparameters for both the baseline and the proposed RL algorithms are in accordance with (Zhou and Gayah, 2021) and summarized in Table 2. Note that the minimal learning rate values are relatively larger than common RL algorithms, since we aim to get a trade-off between optimality and adaptiveness when the algorithm is experiencing disruptions. The number of simulations generated under the Ape-X architecture is $32$ per training episode. The coefficients $\xi_{1}$ and $\xi_{2}$ for the damping term $r_{dam,k}$ in Eq. 10 to counter the potential oscillatory actions are set to be $1$ and $6$. The weights $\omega_{h}$ and $\omega_{\Delta h}$ for the redundancy term $r_{red,k}$ is set to be $0.01$ and $0.02$. Both the prediction horizon and control horizon of MPC and MPC-MHE are set to $10$.
Table 2: List of hyperparameters.
@@ 5.2 Performance evaluation
To showcase antifragility and the associated performance loss, as represented by the shaded area in Fig. 12, the main performance indicator being evaluated in this work is the TTS, which is computed by adding up the number of vehicles within the network at each second of the simulation. Although the reward in the RL algorithm is defined based on trip completion with additional damping and redundancy terms, two scenarios with the same trip completion at the end of the simulation may exhibit distinct Total Time Spent (TTS), and the one with the lower value should be regarded as having demonstrated a superior overall performance.
As Sun et al. (2024) has recently proved the fragile nature of urban road networks, fully antifragile traffic control strategies may be unattainable. When the magnitude of disruptions is sufficiently large, urban road networks will inevitably become fragile due to their real-world constraints. Therefore, this paper aims to demonstrate that the proposed perimeter control algorithm is relatively more antifragile than the state-of-the-art baseline algorithms. To quantify the antifragility of different algorithms, we calculate the distribution skewness based on the samples from the last $N_{episode}=25$ incremental episodes, with $\mu$ and $\sigma$ denoting their mean and the standard deviation:
@@
A skewness computed as 0 indicates that the system itself or the applied algorithm makes the system neither fragile nor antifragile. Negative skewness indicates the distribution has a longer or fatter left tail and thus a higher degree of concavity in the performance function, which showcases antifragility. On the other hand, a positive skewness signifies fragility. For the detailed discussion regarding skewness and antifragility, interested readers may refer to Sun et al. (2024). Therefore, in addition to demonstrating a superior performance in terms of lower TTS, the proposed method should also yield a lower skewness than the benchmarks.
@@ 6.1 Performance under full observability
Under idealized full observability, all algorithms are given perfect OD information $n_{ij}(t),q_{ij}(t)]$ without disruptions. As discussed in Section 5, MPC-MHE is only applied under supply disruption, where a linearly increasing reduction factor $r$ is assumed and can be estimated through MHE.
@@ 6.1.1 Demand disruption
Demand disruption can be understood as a surging traffic demand within a short period, as illustrated in Fig. 7(a) on top of $q_{22,G}$ as a critical scenario. With Fig. LABEL:fig:_incr_d, we present the performance curves of different studied algorithms under the cordon-shaped perimeter control environment, i.e., MPC as in Geroliminis et al. (2013) in orange, baseline RL in Zhou and Gayah (2021) in gray, and the proposed antifragile RL algorithm in blue. The first 50 episodes represent the training process for the RL-based algorithms under the static demand profile in Fig. 7(a). The TTS curves drop down and reach a comparable performance as MPC, corresponding to the conclusion in Zhou and Gayah (2021). The 50 episodes are then followed by another 25 episodes with linearly increasing demand disruption, with no stochasticity assumed on top of the linearly growing magnitude. The studied algorithms demonstrate distinct capabilities of learning and adapting to disruptions and performance variance. The proposed antifragile RL algorithm exhibits both superior performance and reduced variance, indicated by its significantly narrower shaded area. Furthermore, the performance curve of the proposed method also appears less convex than the other algorithms, demonstrating its relative antifragility. Note that due to the inherent variability of RL-based algorithms, as well as the high sensitivity of the distribution skewness on such variability, which will be computed below, we apply a sliding window of 5 episodes to smooth the performance curve.
@@
Such superior performance between episode 507550-75 under incremental demand is further quantitatively gauged with a polar plot in Fig. LABEL:fig:_incr_d_p, which shows the performance difference in TTS reduction normalized over the baseline RL algorithm. Instead of variance, the shaded blue area highlights the performance gain of our proposed method compared to the baseline RL algorithm, showing a steady increase in TTS reduction and being largely concave. The average performance gain is $9.3\%$ and reaches $21.0\%$ after 25 episodes under incremental demand disruption. On the other hand, although MPC performs relatively better when the magnitude of disruption is low, its performance quickly draws near to the baseline RL under high demand disruptions, achieving only a $4.8\%$ improvement at peak disruption.
@@
To quantify the antifragile properties of each perimeter control algorithm, we compute the distribution skewness for each method up to the $n$-th episode, as in Fig. LABEL:fig:_incr_d_sk. The value of the skewness curve at the final episode reflects the antifragility characteristics across all $25$ episodes under incremental disruptions. Since the calculation of skewness requires a minimum sample size to yield meaningful results, a buffer of $5$ episodes is included at the start of the incremental disruption phase. The skewness of the baseline RL in gray rises rapidly, while the skewness of the MPC algorithm progresses even faster to $0.86$. On the contrary, the skewness curve in blue of the antifragile RL-based algorithm increases more slowly and remains largely below the baseline curve, ultimately achieving the lowest final skewness at $0.46$.
@@
Fig. LABEL:fig:_d_un shows the performance when stochasticity is introduced into the magnitude of disruptions, following a disruption shuffling scheme discussed in Section 5.1. The results display a consistent trend with the non-stochastic scenario, but with the performance curves being less smooth, as in Fig. LABEL:fig:_incr_d_p_un, suggesting that the presence of uncertainty adds complexity to the training process. Although the performance of both RL-based methods deteriorates slightly, the proposed antifragile RL algorithm still outperforms the baseline RL by an average of $8.0\%$ and achieves $18.6\%$ performance gain at the end of the simulation under incremental demand disruptions, as well as a gain of $12.9\%$ compared to MPC. The fact that the TTS difference curve is approximately convex for MPC and concave for the antifragile RL algorithm can be quantified in Fig. LABEL:fig:_incr_d_sk_un. While the final distribution skewness of the baseline and the antifragile RL algorithms is $0.70$ and $0.56$, respectively, the skewness of MPC records a value of $0.93$, indicating its high fragility under incremental demand disruptions. An interesting phenomenon worth noticing is the spike of the skewness curve of the proposed algorithm around episode 606560-65 in LABEL:fig:_incr_d_sk_un, which can also be observed in Fig. LABEL:fig:_incr_d_sk. The spike is likely due to the interplay between multiple RL hyperparameters. Particularly after episode $50$, the replay buffer contains a mixture of training data under both disrupted and undisrupted scenarios, resulting in learning from a mixed dataset and causing a moderate deviation from optimal performance.
As the major contribution of the present research is the design of the antifragile module, which can be plugged onto other developed RL algorithms, enabling such algorithms to perform superior and become antifragile when faced with disruptions of increasing levels. Therefore, most hyperparameters of the proposed algorithm with the antifragile module should largely align with the baseline algorithm. We thus implement a sensitivity analysis on the reward weights, which is essential as a methodological contribution, and learning rates, which deviates from the original work in Zhou and Gayah (2021) due to experimental design. Table 3 summarizes the TTS at the peak disruption magnitude and the maximal skewness. We tested different magnitudes of the weight parameters in the redundancy term and found $\omega_{h}=0.01$ and $\omega_{\Delta h}=0.02$ yield the best results with both the lowest TTS and skewness. And as reward shaping heavily relies on a harmony of various reward terms, the RL agent’s decision-making can also be compromised when the weight parameters are too high. For the learning rate of the actor $lr_{a}$ and the critic $lr_{c}$, we proportionally changed the initial learning rates. Results in Table 3 demonstrate that $lr_{a}=0.008$ and $lr_{c}=0.004$ offer the best performance. Smaller learning rates will diminish the learning capability of the agent, especially under increasing disruptions. When the learning rates are high, for instance, $lr_{a}=0.016$ and $lr_{c}=0.008$, although the performance and skewness are comparable or even better than the optimal learning rate, there is a likelihood of causing gradient explosion. Likewise, $lr_{a}=0.032$ and $lr_{c}=0.016$ lead to more frequent exploding updates. For training stability, $lr_{a}=0.008$ and $lr_{c}=0.004$ are selected as the final learning rates.
Table 3: Sensitivity analysis on the weight parameters and the initial learning rates.
Weight parameters Performance Skewness Learning rate Performance Skewness
$\omega_{h}=0.000,\omega_{\Delta h}=0.000$ $1.085\cdot 10^{8}$ $0.55$ $lr_{a}=0.002,lr_{c}=0.001$ $2.282\cdot 10^{8}$ $0.92$
$\omega_{h}=0.001,\omega_{\Delta h}=0.002$ $1.048\cdot 10^{8}$ $0.49$ $lr_{a}=0.004,lr_{c}=0.002$ $1.923\cdot 10^{8}$ $0.80$
$\omega_{h}=0.010,\omega_{\Delta h}=0.020$ $1.026\cdot 10^{8}$ $0.46$ $lr_{a}=0.008,lr_{c}=0.004$ $1.026\cdot 10^{8}$ $0.46$
$\omega_{h}=0.100,\omega_{\Delta h}=0.200$ $1.217\cdot 10^{8}$ $0.62$ $lr_{a}=0.016,lr_{c}=0.008$ $1.008\cdot 10^{8}$ $0.48$
Perimeter control is based on MFD and is therefore intrinsically computationally efficient. Zhou and Gayah (2021) discussed the computation costs and highlighted that although the training time for the RL agent is higher than MPC, the computation power required for the testing phase is lower. In our experiment, both RL-based algorithms and MPC require only seconds to compute the perimeter control variables. However, as the proposed antifragile RL algorithm incorporates a higher state dimension for calculating the derivatives compared to the baseline algorithm, we further study the computation time difference for training the neural networks between these two RL-based algorithms, showing a moderate $18.7\%$ increase. As the neural network for the baseline algorithm is a relatively simple two-layer network but sufficient for solving the perimeter control problem, the training can also be achieved within seconds, and the $18.7\%$ increase is essentially negligible.